• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name silwal

    Log entry time 14:48:00 on November 11, 2009

    Entry number 298996


    keyword=dithering, still not working

    The attached plots below were taken on 10Nov09, swing shift. The plots
    below show that the SL20 magnet kick (pink trace) is working fine, and is
    of the same size as that during HAPPEX III. But the other coils fail to
    produce much beam displacement, not sure why yet.

    The EPICS readback on the magnet EPICS gui is working fine; we can see
    the coil amplitude change the way it is supposed to during dithering cycle.

    Figure 1

    Figure 2